# Comparing Targeting Strategies for Maximizing Social Welfare with Limited Resources

### v1

@@ Abstract
Machine learning is increasingly used to select which individuals receive limited-resource interventions in domains such as human services, education, development, and more. However, it is often not apparent what the right quantity is for models to predict. In particular, policymakers rarely have access to data from a randomized controlled trial (RCT) that would enable accurate estimates of treatment effects – which individuals would benefit more from the intervention. Observational data is more likely to be available, creating a substantial risk of bias in treatment effect estimates. Practitioners instead commonly use a technique termed "risk-based targeting" where the model is just used to predict each individual’s status quo outcome (an easier, non-causal task). Those with higher predicted risk are offered treatment. There is currently almost no empirical evidence to inform which choices lead to the most effect machine learning-informed targeting strategies in social domains. In this work, we use data from 5 real-world RCTs in a variety of domains to empirically assess such choices. We find that risk-based targeting is almost always inferior to targeting based on even biased estimates of treatment effects. Moreover, these results hold even when the policymaker has strong normative preferences for assisting higher-risk individuals. Our results imply that, despite the widespread use of risk prediction models in applied settings, practitioners may be better off incorporating even weak evidence about heterogeneous causal effects to inform targeting.
@@ 1 Introduction
The main challenge is that heterogeneous treatment effects are difficult to learn: doing so requires a potentially large amount of data that is unconfounded, i.e., where treatment is assigned in a manner (conditionally) independent of each individual’s outcomes. In an idealized setting, one could conduct a randomized controlled trial (RCT) and experimentally find the subpopulations that benefit most from a treatment. However, this is not always possible for two reasons: 1) Conducting an RCT takes time (potentially years) and resources that the policymaker may not be willing to spend. 2) In some domains, there are ethical objections to experimentation. For example, while the RCT is being conducted, people in genuine need of the treatment could be assigned to the control group and suffer negative outcomes. Policymakers may prefer to prioritize access to treatment via an already-available proxy metric that is believed to align with need if experimenting to gather additional evidence would be seen as unethical.
@@
One particularly common proxy is to target according to the baseline risk each individual faces, i.e., their expected outcome in the absence of treatment (as opposed to the treatment effect, which is the difference in outcomes induced by treatment). This strategy has been referred to as risk-based targeting [38]. Individuals with poor predicted baseline outcomes may be seen as needing assistance the most. Policymakers may also believe that these individuals also stand the most to benefit since their status-quo prognosis is the worst. Importantly, baseline risks can often be learned using existing administrative data (from before a treatment was introduced) instead of requiring a new experiment, making this strategy easily implementable in many practical settings. For all of these reasons, risk-based targeting has seen widespread use by policymakers in a wide range of domains, including targeting humanitarian assistance [2], allocating homelessness services via vulnerability scores [32], and the use of “early warning systems" in education [29].
@@
Despite this widespread usage, there is only a limited amount of work which empirically assesses the effectiveness of risk-based targeting: do individuals with the greatest baseline risk actually tend to benefit the most from intervention? The few existing studies speak only to a single, specific application domain each. [4] study an RCT where students in a university program were provided a nudge (treatment) as a reminder to renew their financial aid application, concluding that students with intermediate non-renewal risk saw the largest treatment effect. Students with greatest risk of non-renewal, who would be targeted under a risk-based strategy, saw less benefit. [3] study a marketing domain and use two field experiments to show that targeting high-risk customers, or customers likely to churn as predicted by a machine learning model, can be ineffective, encouraging practitioners to use RCTs to better inform their decision. However, as discussed above, running a RCT may be infeasible in many settings. The alternative more likely to be available to practitioners is to simply estimate treatment effects using observational data which likely suffers from confounding, potentially leading to biased estimates of treatment effects.
@@
First, we assess the efficacy of risk-based targeting on a wider variety of real RCT datasets encompassing settings in economics, healthcare and education in contrast to prior studies that generally focus on one dataset. We demonstrate how high values of baseline risk do not always correspond to high treatment effects, rendering risk-based targeting suboptimal if the policymaker adopts a typical utilitarian goal of maximizing the expected improvement in outcomes from the intervention.
@@
Second, we compare risk-based targeting to targeting policies based on biased estimates of treatment effect obtained from confounded data. Such biased estimates are likely when conducting a full-fledged RCT is infeasible and policymakers have to rely on available observational data alone. Accordingly, potentially-based causal estimates represent the likely alternative to risk-based targeting in many domains of practical interest. To our knowledge, these two strategies have not been explicitly compared. We find that across even relatively severe levels of confounding, a utilitarian policymaker often prefers targeting according to biased estimates of the treatment effect rather than baseline risk.
@@
Finally, we analyze the setting where a policymaker has inequality-averse preferences: oftentimes, policymakers may prefer interventions which benefit those who are worse-off to begin with even if they produce less aggregate impact. Such normative goals are one possible justification for risk-based targeting, even if risk-based targeting is less attractive in standard utilitarian terms. We compare the two targeting strategies under popular classes of social welfare functions which capture inequality-averse preferences. We find that treatment effect based targeting is almost always favorable to risk-based targeting, even in scenarios where policy makers are inequality-averse and the only available data is confounded to some degree.
2 Related Work
@@
Measuring heterogeneity in treatment effect and choosing which subpopulations to assign a treatment to has long been an active avenue of research in causal inference literature with a variety of methods proposed to solve this problem. [13, 15, 14, 11, 37] use forest-based algorithms to identify groups that show heterogeneity in treatment effect with other identified groups. [35] proposed to measure the interaction between treatment and covariates by numerically binarizing the treatment and including the products of this variable with each covariate in a regression model. [22] uses meta-learners that decompose estimating the CATE into several sub-regression problems that can be solved with any regression or supervised learning method. The problem of choosing who to treat is closely related to identifying the heterogeneity in treatment effects. This often involves balancing policies based solely on estimates of conditional average treatment effect (CATE) with additional prioritization rules set by the policymaker. [39] proposes rank-weighted average treatment effect metrics for testing the quality of treatment prioritization rules, providing an example involving optimal targeting of aspirin to stroke patients.
@@ 3 Preliminaries
Consider a setting where there is a population of individuals who are candidates for a treatment or intervention. Each individual has a feature vector $X\in\mathbb{R}^{d}$. Here we are concerned with binary treatments. Following Neyman ([34]) and Ruben’s ([31]) potential outcomes framework, we use $Y^{(1)}$ to denote the outcome that an individual would experience under treatment and $Y^{(0)}$ to denote the outcome they would experience if not treated. Their individual treatment effect, quantifying their benefit from receiving treatment, is $Y^{(1)}-Y^{(0)}$. We assume that $(X,Y^{(0)},Y^{(1)})$ are drawn i.i.d. for each individual from some joint distribution. In order to identify individuals who are likely to benefit, a common strategy is to use individuals’ observed covariates to predict the expected treatment effect. The conditional average treatment effect (CATE) at $X=x$ is defined as:
@@
Estimating the CATE in order to target based on treatment effects is a difficult statistical problem. Suppose we have access to data corresponding to $n$ people, labeled $i=1,...,n$, consisting of features $X_{i}$, a treatment assignment $W_{i}\in\{0,1\}$, and the observed outcome $Y_{i}=Y_{i}^{(W_{i})}$. Importantly, for each individual, we can observe only the outcome corresponding to the treatment they were actually assigned. Accordingly, identifying treatment effects typically requires a a no-unobserved-confounders assumption [30]:
$\{Y^{(0)},Y^{(1)}\}\perp\!\!\!\perp W\mid X.$ (2)
This assumption is most credible in the context of a randomized controlled trial (RCT). In an RCT, the assignment of treatment, represented by $W_{i}$, is assigned independently of the potential outcomes $Y_{i}$ (potentially after stratification on covariates $X_{i}$). When data is purely observational, practitioners typically try to select a sufficiently rich set of covariates $X$ such that all potential confounders between outcomes and treatment assignment are measured. However, ensuring that all confounders are completely captured is notoriously difficult in practice, creating the likelihood that some bias in the estimated CATE remains [23, 28, 33, 25].
@@
The example from [4] makes the distinction between these two methods clearer. In an experiment where the objective was to "nudge" or remind students in a college program to renew their financial aid applications, targeting based on baseline risk assumes that students predicted to be least likely to renew their aid applications (as determined by a machine learning model) should be prioritized. Meanwhile, the results of the RCT conducted during this experiment show that high values of treatment effect correspond to students with intermediate likelihoods of renewing their aid applications prior to treatment, demonstrating in a disparity in the two methods of targeting.
@@ 4 Methods
4.1   Overview
Our goal is to compare risk-based targeting to treatment effect-based targeting on possibly confounded datasets, with varying degrees of confounding and under different social welfare functions for the policymakers making the treatment assignment polciy. To enable this comparison, we use a range of datasets from real-world randomized controlled trials (RCTs). Using RCTs enables us to credibly estimate heterogeneous treatment effects since the no-unobserved-confounders assumption is guaranteed to be satisfied. Then, we simulate each of the targeting policies of interest and compare their effectiveness with respect to the randomization-enabled treatment effect estimates, under varying utility functions for the policymaker. We now detail the methodology used in each step of this process, starting with the datasets used.
@@ 4.2   Datasets
  • Acupuncture Dataset: This RCT aimed to determine the effect of acupuncture therapy on headache severity in patients with chronic headaches. Our outcome variable is headache severity 1 year post-randomization. Here again, the effect of the treatment is to reduce the severity of headaches, hence the treatment effect is negative. In order to maintain consistency with other plots, we present results with the sign for treatment effect reversed.
  • Tennessee’s Student Teacher Achievement Ratio (STAR) project [1]: The Tennessee State Department of Education conducted a comprehensive four-year study called the Student/Teacher Achievement Ratio (STAR) to examine the effects of class size on student performance. The study design included three different classroom configurations: 1) Small classes with 13-17 students per teacher, 2) Regular classes with 22-25 students per teacher, 3) Regular classes with 22-25 students plus a full-time teacher’s aide. In this paper, we only focus on the first two types of classes mentioned above, so as to maintain consistency with treatment value being binary in other RCTs. We focus on students in kindergarten and a cumulative measure of their scores on various tests as the outcome under consideration.
For each of these datasets, we estimate $E[Y^{(0)}|X]$ using a machine learning model applied to the RCT’s control group and set $b(X)=E[Y^{(0)}|X]$ or $b(X)=-E[Y^{(0)}|X]$ as appropriate (i.e., depending on whether larger outcome values are better or worse). For instance, children with lower baseline test scores in the STAR dataset and patients with high baseline headache severity in the Acupuncture Dataset are considered to be high risk. Additional details about preprocessing for all datasets are included in A.
@@ 4.3   Estimating heterogeneous treatment effects
4.4   Exploring treatment effect heterogeneity with respect to baseline risk
Our first analysis tests one potential rationale for risk-based targeting strategies: the hypothesis that individuals with greater baseline risk will also tend to have greater treatment effects. We frame this as estimating $E[Y^{(1)}-Y^{(0)}|b(X)]$, a conditional average treatment effect just with respect to value of the risk score $b$.
We follow the doubly-robust approach to estimating CATEs, where the pseudo-outcome difference $\chi_{i}(1)-\chi_{i}(0)$ is regressed on the covariates of interest [18]. Because our covariate of interest, $b$, is one-dimensional, we use a kernel regression method to estimate the CATE as a generic smooth function. Specifically, we employ a Gaussian kernel smoothing method. We sort every individual/household by their baseline risk and center an adaptive Gaussian kernel about each data point. The bandwidth of each kernel is determined adaptively based on the local density of the data, defined as half the range of baseline risk values within a fixed window of 200 data points. This ensures that we are able to estimate greater variation in data-rich regions of the space, while imposing greater smoothness at the extremes where less data is present.
$\sigma_{i}=\frac{1}{2}(b(X_{i+100})-b(X_{i-99}))$ (4)
$\text{CI}=\hat{\tau}(b(X_{i}))\pm 1.96\sqrt{\frac{\sum_{j=1}^{n}K(\frac{b(X_{j})-b(X_{i})}{\sigma_{i}})^{2}(\hat{\tau}_{j}-\hat{\tau}(b(X_{i})))^{2}}{(\sum_{j=1}^{n}K(\frac{b(X_{j})-b(X_{i})}{\sigma_{i}}))^{2}}}$ (5)
@@ 4.5   Introducing Confounding
We introduce varying levels of confounding to the RCTs that we study. We do this by simulating adverse selection into treatment, where units are more likely to be observed if the estimated individual-level treatment effects deviate from the mean. Specifically, we generate the biased “observational" dataset by removing data in a systematic manner. This process, inspired by [17], is controlled by a parameter $k$ giving the fraction of data removed, with higher $k$ corresponding to more biased estimates.
From the treated units, we remove the examples that lie in the top $k\%$ percent when ordered in descending order of $(\chi_{i}(1)-\chi_{i}(0))$ (assuming treatment effect is positive) while for the untreated units, we remove the examples that lie in the bottom $k\%$ of examples when ordered in descending order of $(\chi_{i}(1)-\chi_{i}(0))$. In simpler terms, for treated units, we remove examples for which the treatment ’went well’(most positive), while for untreated units, we remove examples for which the lack of treatment did not go well(least positive). The new dataset we obtain now has $k\%$ of the original examples systematically removed to increase confounding and can only provide biased estimates of treatment effect.
Our goal is to compare policies learned using this biased data to risk-based targeting. This evaluation is enabled by access to the original, randomized data. Consider a hypothetical policy that assigns treatments $A(X)\in\{0,1\}$ as a function of individuals’ features $X$. The mean outcome under policy $A$, $E[Y^{(A(X))}]$, can be decomposed as
$E[Y^{(A(X))]}=E[Y^{(0)}]+Pr(A(X)=1)E[Y^{(1)}-Y^{(0)}|A(X)=1].$
The term $E[Y^{(1)}-Y^{(0)}|A(X)=1]$ represents the treatment effect on the treated population and captures how much the policy improves over no treatment. For policies with the same budget (equal $\Pr(A(X)=1)$), only this term varies and so we assess allocation policies by their expected treatment-on-the-treated. Following standard doubly-robust estimators for (group) average treatment effects [19], we empirically estimate this quantity as
$\frac{1}{\sum_{j=1}^{n}A(X_{j})}\sum_{j=1}^{n}A(X_{j})(\chi_{j}(1)-\chi_{j}(0)).$

### v2

@@ Abstract
Machine learning is increasingly used to select which individuals receive limited-resource interventions in domains such as human services, education, development, and more. However, it is often not apparent what the right quantity is for models to predict. Policymakers rarely have access to data from a randomized controlled trial (RCT) that would enable accurate estimates of which individuals would benefit more from the intervention, while observational data creates a substantial risk of bias in treatment effect estimates. Practitioners instead commonly use a technique termed “risk-based targeting" where the model is just used to predict each individual’s status quo outcome (an easier, non-causal task). Those with higher predicted risk are offered treatment. There is currently almost no empirical evidence to inform which choices lead to the most effective machine learning-informed targeting strategies in social domains. In this work, we use data from 5 real-world RCTs in a variety of domains to empirically assess such choices. We find that when treatment effects can be estimated with high accuracy (which we simulate by allowing the model to partially observe outcomes in advance), treatment effect based targeting substantially outperforms risk-based targeting, even when treatment effect estimates are biased. Moreover, these results hold even when the policymaker has strong normative preferences for assisting higher-risk individuals. However, the features and data actually available in most RCTs we examine do not suffice for accurate estimates of heterogeneous treatment effects. Our results suggest treatment effect targeting has significant potential benefits, but realizing these benefits requires improvements to data collection and model training beyond what is currently common in practice.
@@ 1 Introduction
The main challenge is that heterogeneous treatment effects are difficult to learn: doing so requires a potentially large amount of data that is unconfounded, i.e., where treatment is assigned in a manner (conditionally) independent of each individual’s outcomes. Conducting a randomized controlled trial (RCT) to achieve this is often infeasible due to time constraints, resource limitations, and ethical concerns. Policymakers may prefer to prioritize access to treatment via an already-available proxy metric that is believed to align with need if experimenting to gather additional evidence would be seen as unethical.
@@
One particularly common proxy is to target according to the baseline risk each individual faces, i.e., their expected outcome in the absence of treatment (as opposed to the treatment effect, which is the difference in outcomes induced by treatment). This strategy has been referred to as risk-based targeting [42]. Individuals with poor predicted baseline outcomes may be seen as needing assistance the most. Policymakers may also believe that these individuals also stand the most to benefit since their status-quo prognosis is the worst. Importantly, baseline risks can often be learned using existing administrative data (from before a treatment was introduced) instead of requiring a new experiment, making this strategy easily implementable in many practical settings. For all of these reasons, risk-based targeting has seen widespread use by policymakers in a wide range of domains, including targeting humanitarian assistance [2], allocating homelessness services via vulnerability scores [35], and the use of "early warning systems" in education [32].
@@
Despite this widespread usage, there is only a limited amount of work which empirically assesses the effectiveness of risk-based targeting: do individuals with the greatest baseline risk actually tend to benefit the most from intervention? Existing studies speak only to a single specific application domain each, for instance, targeting students for sending reminders for financial aid applications [4], preventing customer churn [3], or targeting cash transfers [15]. [23] examines this from a human perspective by empirically eliciting decision-maker preferences for whether to prioritize more vulnerable households or households who would best take advantage of more intensive interventions when determining which homeless households to serve with limited housing assistance. [3] encourages practitioners to use RCTs to better inform their decision, however, as discussed above, running a RCT may be infeasible in many settings. The alternative more likely to be available to practitioners is to simply estimate treatment effects using observational data which likely suffers from confounding, potentially leading to biased estimates of treatment effects.
@@
First, we assess the efficacy of risk-based targeting on a wider variety of real RCT datasets encompassing settings in economics, healthcare and education in contrast to prior studies that generally focus on one dataset. We find a generally noisy and variable relationship between baseline risk and treatment effects: high-risk individuals seem to benefit more on average in most domains, but with substantial variance in treatment effects which is not explained by baseline risk. When treatment effects can be estimated reliably (which we simulate by allowing the model to partially observe outcomes in advance), targeting based on these estimates produces substantially better results compared to risk-based targeting. However, in practice, treatment effects must be predicted from features alone, and limited data can make it difficult to learn accurate mappings from features to treatment effects, potentially degrading the performance of treatment effect based targeting.
@@
Second, we compare risk-based targeting to targeting policies based on biased estimates of treatment effect obtained from confounded data. Such biased estimates are likely when conducting a full-fledged RCT is infeasible and policymakers have to rely on available observational data alone. Accordingly, potentially-biased causal estimates represent the likely alternative to risk-based targeting in many domains of practical interest. To our knowledge, these two strategies have not been explicitly compared. We find that across even relatively severe levels of confounding, a utilitarian policymaker often prefers targeting according to biased estimates of the treatment effect rather than baseline risk when treatment effects can be estimated reliably at no confounding.
@@
Finally, we analyze the setting where a policymaker has inequality-averse preferences: oftentimes, policymakers may prefer interventions which benefit those who are worse-off to begin with even if they produce less aggregate impact. Such normative goals are one possible justification for risk-based targeting, even if risk-based targeting is less attractive in standard utilitarian terms. We compare the two targeting strategies under popular classes of social welfare functions which capture inequality-averse preferences. We find that when treatment effects can be estimated reliably, treatment effect based targeting is typically favorable to risk-based targeting, even in scenarios where policy makers are inequality-averse and the only available data is confounded to some degree.
All together, our results suggest that the barrier to targeting on treatment effects is more likely to lie in the collection of large and rich enough datasets to support accurate prediction of outcomes, while confounding or inequality-averse preferences have more limited impact. This implies that large-scale administrative datasets may be a more attractive avenue for policy learning than previously thought, even if they lack the guaranteed validity of a randomized controlled trial.
@@
Related Work: Measuring heterogeneity in treatment effect and choosing which subpopulations to assign a treatment to has long been an active avenue of research in causal inference literature with a variety of methods proposed to solve this problem. [13, 17, 16, 11, 41] use forest-based algorithms to identify groups that show heterogeneity in treatment effect with other identified groups. [38] proposed to measure the interaction between treatment and covariates by numerically binarizing the treatment and including the products of this variable with each covariate in a regression model. [25] uses meta-learners that decompose estimating the CATE into several sub-regression problems that can be solved with any regression or supervised learning method. The problem of choosing who to treat is closely related to identifying the heterogeneity in treatment effects. This often involves balancing policies based solely on estimates of conditional average treatment effect (CATE) with additional prioritization rules set by the policymaker. [43] proposes rank-weighted average treatment effect metrics for testing the quality of treatment prioritization rules, providing an example involving optimal targeting of aspirin to stroke patients. Meanwhile, [14] provides a theoretical analysis of risk-based targeting strategies, establishing performance bounds for scoring rules when treatments are benign.
@@ 2 Preliminaries
Consider a setting where there is a population of individuals who are candidates for a treatment or intervention. Each individual has a feature vector $X\in\mathbb{R}^{d}$. Here we are concerned with binary treatments. Following Neyman and Ruben’s [37, 34] potential outcomes framework, we use $Y^{(1)}$ to denote the outcome that an individual would experience under treatment and $Y^{(0)}$ to denote the outcome they would experience if not treated. Their individual treatment effect is $Y^{(1)}-Y^{(0)}$. We assume that $(X,Y^{(0)},Y^{(1)})$ are drawn i.i.d. for each individual from some joint distribution. In order to identify individuals who are likely to benefit, a common strategy is to use individuals’ observed covariates to predict the expected treatment effect. The conditional average treatment effect (CATE) is defined as:
@@
Estimating the CATE in order to target based on treatment effects is a difficult statistical problem. Suppose we have access to data corresponding to $n$ people, labeled $i=1,...,n$, consisting of features $X_{i}$, a treatment assignment $W_{i}\in\{0,1\}$, and the observed outcome $Y_{i}=Y_{i}^{(W_{i})}$. Importantly, for each individual, we can observe only the outcome corresponding to the treatment they were actually assigned. Accordingly, identifying treatment effects typically requires a a no-unobserved-confounders assumption [33]: $\{Y^{(0)},Y^{(1)}\}\perp\!\!\!\perp W\mid X$. This assumption is most credible in the context of a randomized controlled trial (RCT). In an RCT, the assignment of treatment, represented by $W_{i}$, is assigned independently of the potential outcomes $Y_{i}$ (potentially after stratification on covariates $X_{i}$). When data is purely observational, practitioners typically try to select a sufficiently rich set of covariates $X$ such that all potential confounders between outcomes and treatment assignment are measured. However, ensuring that all confounders are completely captured is notoriously difficult in practice, creating the likelihood that some bias in the estimated CATE remains [26, 31, 36, 28].
@@ 3 Methods
We compare risk-based and treatment effect-based targeting strategies using real-world RCTs, which allow credible estimation of heterogeneous treatment effects because of the assumption of no unmeasured confounding. We simulate these targeting strategies under varying degrees of confounding and under different social welfare functions for policymakers, evaluating their effectiveness against randomization-enabled treatment effect estimates. We now detail the methodology used in each step of this process, starting with the datasets used.
@@ 3.1   Datasets
Acupuncture Dataset: [40] This RCT aimed to determine the effect of acupuncture therapy on headache severity in patients with chronic headaches. Our outcome variable is headache severity 1 year post-randomization. Here again, the effect of the treatment is to reduce the severity of headaches, hence the treatment effect is negative. In order to maintain consistency with other plots, we present results with the sign for treatment effect reversed.
Tennessee’s Student Teacher Achievement Ratio (STAR) project [1]: This four-year study by the Tennessee State Department of Education examined class size effects on student performance. The design compared: 1) Small classes (13-17 students per teacher), 2) Regular classes (22-25 students), and 3) Regular classes with a teacher’s aide. We focus only on the first two configurations to maintain binary treatment consistency with other RCTs. Our outcome measure is kindergarten students’ cumulative test scores.
For each of these datasets, we estimate $E[Y^{(0)}|X]$ using a machine learning model applied to the RCT’s control group and set $b(X)=E[Y^{(0)}|X]$ or $b(X)=-E[Y^{(0)}|X]$ as appropriate so that larger values of $b$ indicate worse outcomes. Additional details are included in Appendix A.
@@ 3.2   Estimating heterogeneous treatment effects
3.3   Exploring treatment effect heterogeneity
Our first analysis tests one potential rationale for risk-based targeting strategies: the hypothesis that individuals with greater baseline risk will also tend to have greater treatment effects. We frame this as estimating $E[Y^{(1)}-Y^{(0)}|b(X)]$, a conditional average treatment effect just with respect to value of the risk score $b$. We follow the doubly-robust approach to estimating CATEs, where the pseudo-outcome difference $\chi_{i}(1)-\chi_{i}(0)$ is regressed on the covariates of interest [20]. Because our covariate of interest, $b$, is one-dimensional, we use a kernel regression method to estimate the CATE as a generic smooth function (see Appendix A.3).
In later analyses, we compare the welfare gains of potential targeting policies. Here, we also use a doubly-robust estimator. Consider a hypothetical policy that assigns treatments $A(X)\in\{0,1\}$ as a function of individuals’ features $X$. The mean outcome under policy $A$, $E[Y^{(A(X))}]$, can be decomposed as
$E[Y^{(A(X))}]=E[Y^{(0)}]+Pr(A(X)=1)E[Y^{(1)}-Y^{(0)}|A(X)=1].$
The term $E[Y^{(1)}-Y^{(0)}|A(X)=1]$ represents the treatment effect on the treated population and captures how much the policy improves over no treatment. For policies with the same budget (equal $\Pr(A(X)=1)$), only this term varies and so we assess allocation policies by their expected treatment-on-the-treated. Following standard doubly-robust estimators for (group) average treatment effects [21], we empirically estimate this quantity on a separate test set as
$\frac{1}{\sum_{j=1}^{n}A(X_{j})}\sum_{j=1}^{n}A(X_{j})(\chi_{j}(1)-\chi_{j}(0)).$ (2)
@@ 3.4   Introducing Confounding
We introduce varying levels of confounding to the RCTs that we study. We do this by simulating adverse selection into treatment, where units are more likely to be observed if the estimated individual-level treatment effects deviate from the mean. Specifically, we generate the biased “observational" dataset by removing data in a systematic manner. This process, inspired by [19], is controlled by a parameter $k$ giving the fraction of data removed, with higher $k$ corresponding to more biased estimates. From the treated units, we remove the examples that lie in the top $k\%$ percent when ordered in descending order of $(\chi_{i}(1)-\chi_{i}(0))$ (assuming treatment effect is positive) while for the untreated units, we remove the examples that lie in the bottom $k\%$ of examples when ordered in descending order of $(\chi_{i}(1)-\chi_{i}(0))$. In simpler terms, for treated units, we remove examples for which the treatment ’went well’(most positive), while for untreated units, we remove examples for which the lack of treatment did not go well(least positive).